# Akaki Mamageishvili Offchain Labs Benny Sudakov ETH Zurich

### v1

@@ Abstract
We compare the efficiency of restaking and Proof-of-Stake (PoS) protocols in terms of stake requirements. First, we consider the sufficient condition for the restaking graph to be secure. We show that the condition implies that it is always possible to transform such a restaking graph into secure PoS protocols. Next, we derive two main results, giving upper and lower bounds on required extra stakes that one needs to add to validators of the secure restaking graph to be able to transform it into secure PoS protocols. In particular, we show that the restaking savings compared to PoS protocols can be very large and can asymptotically grow in the worst case as a square root of the number of validators. We also study a complementary question of transforming secure PoS protocols into an aggregate secure restaking graph and provide lower and upper bounds on the PoS savings compared to restaking.
@@ 1 Introduction
Restaking has recently been proposed as a cryptoeconomically more efficient alternative to the Proof-of-Stake solution concept for blockchains. The latter was proposed on its own as an energy-efficient alternative to the Proof-of-Work concept  [^1] See [6] for a more detailed discussion on Proof-of-Work vs. Proof-of-Stake concepts.. In restaking, validators that are staked for one protocol can reuse their stake to secure other projects or protocols of their choice. This gives a boost to newer projects in attracting capital for their cryptoeconomic security. Cryptoeconomic security refers to a security derived from staked parties being slashed their stakes in case of misbehavior. Restaking seems to be attractive for the validators of the original protocol as well. In particular, restaking allows them to reuse their computational power and storage allocation for additional payoffs coming from other projects’ rewards. The original protocol, for which the validators are staked, may have requirements that do not exhaust all the resources of the validators, usually for the purpose of decentralization [^2] Check out Ethereum network plans to resolve this asymmetry between validator resources using rainbow staking https://ethresear.ch/t/unbundling-staking-towards-rainbow-staking/18683.. Another potential reason for such availability might be economies of scale for establishing further validation services.
@@
The idea of restaking was originally proposed in the whitepaper of the Eigenlayer project, [7]. A similar concept existed in traditional finance literature under the name of rehypothecation[4]. The interdependencies between protocols and validators introduced by restaking create new risks, which have been modeled and analyzed in the original whitepaper, as well as in the series of papers [3][1][8][5]. The simplest way to verify the security of the restaking uses a cost-benefit analysis. In this analysis, the total amount of lost stakes of attackers is compared to the total value obtained from the attack of a set of validators. The same cost-benefit analysis can be used to check the security of a Proof-of-Stake protocol, in which a single project/service is secured by multiple validators. This allows a reasonable comparison of restaking and multiple PoS protocols. [2] discusses ways to estimate the value gained from attacking a protocol and cryptoeconomic security more in detail.
@@
Since the Proof-of-Stake protocol is simpler and hence safer solution concept than restaking, we are interested in checking comparative benefits of restaking protocols. One particular attractive point of restaking protocols can be the lower total staking requirements than that of Proof-of-Stake protocols, since there are opportunity costs associated with locking up stake for security reasons. Most PoS protocols pay for this cost by rewarding stakers with newly issued tokens, which leads to inflation. Hence, minimizing the total stake amount while maintaining security can be a reasonable goal in both restaking or PoS protocols. Starting with a secure restaking protocol and initial stake endowments for all validators, we measure how much extra stake would be required if validators decided to form separate Proof-of-Stake protocols with their stake amounts, dividing them across the same set of projects they secure in the original restaking graph. We only allow all validator stakes to (weakly) increase. This captures a real-life consideration in which validators’ stakes can not be taken away, but their stakes can be increased through financing when needed. If both increasing and decreasing of validator stakes were allowed, it would be trivial to construct a restaking graph with minimum total stake. This amount is equal to the sum of all project values, and secure PoS protocols for all projects can be constructed by dividing the stakes of the validators.
@@
First, we check the sufficient condition for a secure restaking graph, specified by the EigenLayer project [7]. Although this sufficient condition is the only currently available (generic) condition that can be checked in polynomial time in the size of the restaking graph, it does not save on total required stake. In particular, it is possible to divide the initial stakes to obtain secure PoS protocols across all projects. To further extend the analysis, we define restaking savings of a secure restaking graph. It is equal to the extra stake required, relative to the original total available stake in the graph, to be able to obtain secure PoS protocols through the division of stake described above. Higher restaking savings indicate an advantage of restaking to PoS. To upper bound this value, we use simple secure PoS constructions from the original restaking graph. In particular, we upper bound the value of restaking savings with the highest degree among the projects. Another upper bound is the minimum incidence of any validator in the cover of a validator set by project neighborhoods. The third upper bound is equal to second largest multiplicative inverse of the security parameter, which is a small constant value in many practical cases. The last upper bound is equal to the square root of the number of validators. Next, we give a construction to lower bound this value. The lower bound example asymptotically matches all upper bounds obtained.
@@
2 Model
@@
We consider stake vectors $\sigma^{\prime}$ that (weakly) dominate stake vector $\sigma$, that is, $\sigma^{\prime}_{s}\geq\sigma_{s}$ for any $s$. The set of all stake vectors that dominate $\sigma$ is denoted by $D(\sigma)$. Let $T(\sigma):=\sum_{v\in V}\sigma_{v}$. Then, the total extra value to reach from the stake vector $\sigma$ to the stake vector $\sigma^{\prime}$ is $T(\sigma^{\prime})-T(\sigma)$. The rationale behind considering (weakly) dominant vectors of stakes is that we assume that initial stakes cannot be taken away from the validators, but they can be increased if needed.
@@ 3.1 Restaking Savings
There is a minimum total stake requirement to ensure that there is a division in secure PoS protocols. The amount is equal to $\sum_{s\in S}\pi_{s}$. It can be distributed arbitrarily among validators, with the only guarantee that each service can get its own value staked by one of the validators securing it. However, such a distribution does not take into account that validators already have stakes, which cannot be taken away from them. Another consideration is that the restaking graph with the minimum total stake in ${\cal W}(G)$ is not necessarily secure, while we will focus on secure restaking graphs from now on.
@@
Next, we show $RS(G)\leq K$. Consider a cover of vertices in $V$ by neighborhoods of services $s\in R\subseteq S$ that has the minimum maximum incidence number $K$. For each neighborhood $N_{G}(s)$, where $s\in R$, use construction 3: original $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ stakes from the restaking graph to secure service $s$ in the PoS protocol. To be able to do that, we need to add at most $K-1$ times all $\sigma_{v}$ stakes, for any $v\in V$. In this way, we utilize all $\sigma_{v}$ stakes and secure services in the cover set $R$. For service $s$ in $S\setminus R$, we add $\pi_{s}$ stake once to any of its neighbors in $N_{G}(s)$ and use construction $(0,,,0,\pi_{s},0,...,0)$ to secure service $s$ in a corresponding PoS protocol. This makes sure we have secured all services and utilized all stakes of all validators. Then, $RS(G)\leq(K-1)+1=K$. ∎
Note that $RS(G)\leq K$ in particular implies that $RS(G)\leq\max_{s\in S}(d_{G}(s))$, as $K$ is upper bounded by $\max_{s\in S}(d_{G}(s))$. There are examples where $K=\max_{s\in S}(d_{G}(s))$, implying that the upper bound can be as bad as $n$. However, next, we show an upper bound as a function of the number of validators, which is asymptotically much lower than $n$.
@@
We can secure any service $s\in S\setminus S_{L}$ by adding stakes $\pi_{s}$ for to all its neighbors in $V\setminus V_{L}$, that is, by adding $d_{V\setminus V_{L}}(s)\pi_{s}$ extra stakes. By the construction of the set $S_{L}$ we know that $d_{V\setminus V_{L}}(s)$ is upper bounded by $\sqrt{n}$ for any $s\in S\setminus S_{L}$. We can, at the same time, utilize the remaining stakes in $V\setminus V_{L}$, by using a construction of type ($,..,0,\pi_{s}+\sigma_{v},..0)$.0 ) for $v\in V\setminus V_{L}$. The last step is possible, because for each validator in $V\setminus V_{L}$, there must be at least one service in $S\setminus S_{L}$ it secures. The total extra stake is upper bounded by $\sqrt{n}T(\pi_{S\setminus S_{L}})\leq\sqrt{n}T(\sigma)$, where $T(\pi_{S\setminus S_{L}})$ denotes the sum $\sum_{s\in S\setminus S_{L}}\pi_{s}$. By summing up two values of extra stakes used to secure $S_{L}$ and $S\setminus S_{L}$ sets, and also utilizing all stakes in $V$, we get
@@
Note that in the proofs of Proposition 2 and Theorem 1, we did not use any property of $\alpha$ values. Both upper bounds use combinatorial features of the underlying restaking graph $G$. Next, we show an upper bound that is a function of the security parameters $\alpha_{s}$. The upper bound is interesting in practical cases, since the security parameters of most protocols are usually from the set of fractions $\{\frac{1}{4},\frac{1}{3},\frac{1}{2}\}$. These values come from Byzantine fault-tolerant mechanisms, used to achieve consensus on the state of the chain. We show the following upper bound on $RS(G)$:
@@
    set all values in the vector $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ to 00.
@@
Next, we construct an example of a secure restaking graph that derives a lower bound on the restaking savings, that is linear in the number of services and asymptotically matches to all upper bounds derived so far.
@@
For any $m\in\mathbb{N}$, there are instances of a secure restaking graph $G$ in which the restaking savings $RS(G)\in\Theta(m)$.
@@
Assume the number of validators is $n=m^{2}+1$ for some $m\in\mathbb{N}$. The edge set $E$ consists of $m$ edges $(s,m^{2}+1)$, for any $1\leq s\leq m$ and $m^{2}$ edges of type $(s,(s-1)m+j)$, for any $1\leq s\leq m$, $1\leq j\leq m$.
@@
First, we show that $G$ is a secure restaking graph. Note that the large validator, indexed with ${m^{2}+1}$ and with stake $\sigma_{m^{2}+1}=2m$ does not want to participate in any attack, as she is losing at least the total value available across all services. No service $s\in\{1,2,...,m\}$ can be attacked by all validators securing it other than the large validator, since the service security parameter is $\alpha_{s}=\frac{1}{2m+1}$, while all small validators indexed $((s-1)m+1,(s-1)k+2,...,sm)$ make up stake $1$ in total. The total available stakes securing a service $s$ is $12m+$ 1.
@@
    If there is a subset of validators, whose stakes sum up to a number between $1$ and $2$, then it must be that the total stake, $\sum_{i=1}^{m+1}a_{i}$, is at least $12m+$ 1. Otherwise, this subset would be able to attack the service and the attack would be profitable. This implies that the extra stake size is at least $2m=2m+1-m\frac{1}{m}$.
@@
It is easy to show that $O(m^{2})$ total stake is enough to add to validators in $V$, so that the resulting stake vector $\sigma\in\mathcal{W}(G)$. It can, for example, be done by adding the stake $2m^{2}-m$ to the validator $m^{2}+1$ so that it total stake become $(2m+1)m$. Then splitting this stake into $m$ stakes of $12m+$ 1 each in each of $m$ PoS protocols. Small stakers in these protocols can not attack any of the services even if they all coordinate, while the large staker does not find any attack profitable. This proves the claim of the theorem, $RS(G)\in\Theta(m)$.
@@
That is, restaking savings are not bounded by a constant, and they may grow unboundedly as the number of services and validators grow. Note that we need both values to grow unboundedly. Otherwise, upper bounds obtained in Proposition 2 ensure that $RS(G)$ is constant.

### v2

@@ Abstract
We compare the efficiency of secure restaking and Proof-of-Stake (PoS) protocols in terms of stake requirements. First, we consider the sufficient condition for the restaking graph to be secure. We show that the condition implies that it is always possible to transform such a restaking graph into separate secure PoS protocols. Next, we derive two main results, giving upper and lower bounds on required extra stakes that one needs to add to validators of the secure restaking graph to be able to transform it into secure PoS protocols. In particular, we show that the restaking savings compared to PoS protocols can be very large and can asymptotically grow in the worst case as a square root of the number of validators. We also study a complementary question of aggregating secure PoS protocols into a secure restaking graph and provide lower and upper bounds on the PoS savings compared to restaking.
@@ 1 Introduction
Economic security of blockchains evolves over time. Initially, Proof-of-Work was proposed for Bitcoin security, [6]. Soon after that, Proof-of-Stake was proposed as an energy-efficient alternative to the Proof-of-Work concept  [^1] See [7] for a more detailed discussion on Proof-of-Work vs. Proof-of-Stake concepts.. More recently, Restaking has been proposed as a cryptoeconomically more efficient alternative to the Proof-of-Stake solution concept for blockchains. Restaking refers to a setting in which validators that are staked for one protocol can reuse their stake to secure other projects or protocols of their choice. This creates a bipartite graph structure between the projects and the validators.
@@
It can be argued that restaking gives a boost to newer projects in attracting capital for their cryptoeconomic security. Cryptoeconomic security refers to a security derived from staked parties being slashed their stakes in case of misbehavior. Restaking seems to be attractive for the validators of the original protocol as well. In particular, restaking allows them to reuse their computational power and storage allocation for additional payoffs coming from other projects’ rewards. The original protocol, for which the validators are staked, may have requirements that do not exhaust all the resources of the validators, usually for the purpose of decentralization [^2] Check out Ethereum network plans to resolve this asymmetry between validator resources using rainbow staking https://ethresear.ch/t/unbundling-staking-towards-rainbow-staking/18683.. Another potential reason for such availability might be economies of scale for establishing further validation services.
@@
The simplest way to verify the security of the restaking uses a trivial cost-benefit analysis. In this analysis, the total amount of lost stakes of attackers is compared to the total value obtained from the attack of a set of validators. The same cost-benefit analysis can be used to check the security of a Proof-of-Stake protocol, in which a single project/service is secured by multiple validators. This allows a reasonable comparison of restaking and multiple PoS protocols. [2] discusses ways to estimate the value gained from attacking a protocol and cryptoeconomic security more in detail.
Since the Proof-of-Stake protocol is simpler and hence a more natural solution concept than restaking, we are interested in checking comparative benefits of restaking protocols. One particular attractive point of restaking protocols can be the lower total staking requirements than that of Proof-of-Stake protocols, since there are opportunity costs associated with locking up stake for security reasons.
@@
Most PoS protocols pay for this cost by rewarding stakers with newly issued tokens, which leads to inflation. Hence, minimizing the total stake amount while maintaining security can be a reasonable goal in both restaking or PoS protocols. Starting with a secure restaking protocol and initial stake endowments for all validators, we measure how much extra stake would be required if validators decided to form separate Proof-of-Stake protocols with their stake amounts, dividing them across the same set of projects they secure in the original restaking graph.
In the following, we describe an example that gives intuition of the secure restaking graph and its transformation to secure PoS protocols, see the Figure 1.
Example 1.
There are two services, with valuations $1$ and $3$ and security parameters $\alpha$, which is any number in the interval $(0,1)$ and $\frac{1}{2}$. There are two validators, with stakes $2$ and $2$.
The security parameter of a service establishes what share of the total stake securing the service is needed to successfully attack this service. Let us first check that the restaking graph is secure. It is easy to see:
  • the first validator alone can attack only the first service, and it is not profitable;
  • the second validator alone can not attack any service: it is not securing th first service, while the second service has large enough security parameter.
  • and, finally, both validators can attack both services but they lose the same amount of stake as they gain the value.
Next, we check if the validators can distribute their stakes so that both services are secured in the corresponding PoS protocols. First, note that it is the first validator that divides its stake between two services. The first service needs at least the stake amount $1$ from the first validator. The second service is not secure in the PoS protocol with stakes $1$ and $2$. In fact, it can be easily seen that either the first validator needs to have at least a matching stake $2$, or the second validator needs to have at least a stake $3$. In either case, at least an extra stake $1$ is required to secure both services in PoS protocols.
1322 $S$ $V$
Figure 1: Example construction for $m=3$.
We only allow all validator stakes to (weakly) increase. This captures a real-life consideration in which validators’ stakes can not be taken away, but their stakes can be increased through financing when needed. If both increasing and decreasing of validator stakes were allowed, it would be trivial to construct a restaking graph with minimum total stake. This amount is equal to the sum of all project values, and secure PoS protocols for all projects can be constructed by dividing the stakes of the validators.
@@
  1. 1.
    We check the sufficient condition for a secure restaking graph, specified by the EigenLayer project [8]. This sufficient condition is the only currently available (generic) condition that can be checked in polynomial time in the size of the restaking graph. Restaking graphs that satisfy the sufficient condition do not save on the total required stake. In particular, it is possible to divide the initial stakes of the validators to obtain secure PoS protocols across all projects.
  2. 2.
    To further extend the analysis, we define restaking savings of a secure restaking graph. It is equal to the extra stake required, relative to the original total available stake in the graph, to be able to obtain secure PoS protocols through the division of stake described above. Higher restaking savings indicate an advantage of restaking to PoS.
  3. 3.
    To upper bound restaking savings, we use simple secure PoS constructions from the original restaking graph. In particular, we upper bound the value of restaking savings with the highest degree among the projects. Another upper bound is the minimum incidence of any validator in the cover of a validator set by project neighborhoods. The third upper bound is equal to second largest multiplicative inverse of the security parameter, which is a small constant value in many practical cases. The last upper bound is equal to the square root of the number of validators. Next, we give a construction to lower bound this value. The lower bound example asymptotically matches all upper bounds obtained.
  4. 4.
@@
Related Work
The idea of restaking was originally proposed in the whitepaper of the Eigenlayer project, [8]. A similar concept existed in traditional finance literature under the name of rehypothecation[4]. The interdependencies between protocols and validators introduced by restaking create new risks, which have been modeled and analyzed in the original whitepaper, as well as in a follow up paper [3], in which the authors study robust restaking networks, resistant to a failure of a fraction of validating stakes.  [1] studies a setting in which the value of attacked set of services is not the sum of values in the set, but a concave function. [9] studies elastic restaking services, in which allocated stakes to services make up more than the total stake. [5] studies risks and benefits of combining different restaking services.
2 Notation and Model
@@
We consider stake vectors $\sigma^{\prime}$ that (weakly) dominate stake vector $\sigma$, that is, $\sigma^{\prime}_{v}\geq\sigma_{v}$ for any $v\in V$. The set of all stake vectors that dominate $\sigma$ is denoted by $D(\sigma)$. Let $T(\sigma):=\sum_{v\in V}\sigma_{v}$. Then, the total extra value to reach from the stake vector $\sigma$ to the stake vector $\sigma^{\prime}$ is $T(\sigma^{\prime})-T(\sigma)$. The rationale behind considering (weakly) dominant vectors of stakes is that we assume that initial stakes cannot be taken away from the validators, but they can be increased if needed.
@@ 3.1 Restaking Savings
There is a minimum total stake requirement to ensure that there is a division in secure PoS protocols.
Proposition 2.
The minimum stake amount to guarantee secure PoS protocols is equal to $\sum_{s\in S}\pi_{s}$.
Proof.
The total stake amount $\sum_{s\in S}\pi_{s}$ can be distributed arbitrarily among validators, with the only guarantee that each service can get its own value staked by one of the validators securing it. ∎
However, such a distribution does not take into account that validators already have stakes, which cannot be taken away from them.
Another consideration is that the restaking graph with the minimum total stake in ${\cal W}(G)$ is not necessarily secure, while we will focus on secure restaking graphs from now on.
@@
Next, we show $RS(G)\leq K$. Consider a cover of vertices in $V$ by neighborhoods of services $s\in R\subseteq S$ that has the minimum maximum incidence number $K$. For each neighborhood $N_{G}(s)$, where $s\in R$, use construction 3: original $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ stakes from the restaking graph to secure service $s$ in the PoS protocol. To be able to do that, we need to add at most K1K-1 times all $\sigma_{v}$ stakes, for any $v\in V$. In this way, we utilize all $\sigma_{v}$ stakes and secure services in the cover set $R$. For service $s$ in $S\setminus R$, we add $\pi_{s}$ stake once to any of its neighbors in $N_{G}(s)$ and use construction $(0,,,0,\pi_{s},0,...,0)$ to secure service $s$ in a corresponding PoS protocol. This makes sure we have secured all services and utilized all stakes of all validators. Then, $RS(G)\leq(K-1)+1=K$. ∎
Note that $RS(G)\leq K$ in particular implies that $RS(G)\leq\max_{s\in S}(d_{G}(s))$, as $K$ is upper bounded by $\max_{s\in S}(d_{G}(s))$. There are examples where $K=\max_{s\in S}(d_{G}(s))$, implying that the upper bound can be as high as $n$. However, next, we show an upper bound as a function of the number of validators, which is asymptotically much lower than $n$.
@@
We can secure any service $s\in S\setminus S_{L}$ by adding stakes $\pi_{s}$ for to all its neighbors in $V\setminus V_{L}$, that is, by adding $d_{V\setminus V_{L}}(s)\pi_{s}$ extra stakes. By the construction of the set $S_{L}$ we know that $d_{V\setminus V_{L}}(s)$ is upper bounded by $\sqrt{n}$ for any $s\in S\setminus S_{L}$. We can, at the same time, utilize the remaining stakes in $V\setminus V_{L}$, by using a construction of type $(0,..,0,\pi_{s}+\sigma_{v},..0)$ for $v\in V\setminus V_{L}$. The last step is possible, because for each validator in $V\setminus V_{L}$, there must be at least one service in $S\setminus S_{L}$ it secures. The total extra stake is upper bounded by $\sqrt{n}T(\pi_{S\setminus S_{L}})\leq\sqrt{n}T(\sigma)$, where $T(\pi_{S\setminus S_{L}})$ denotes the sum $\sum_{s\in S\setminus S_{L}}\pi_{s}$. By summing up two values of extra stakes used to secure $S_{L}$ and $S\setminus S_{L}$ sets, and also utilizing all stakes in $V$, we get
@@
Note that in the proofs of Proposition 3 and Theorem 1, we did not use any property of $\alpha$ values. Both upper bounds use combinatorial features of the underlying restaking graph $G$. Next, we show an upper bound that is a function of the security parameters $\alpha_{s}$. We show the following upper bound on $RS(G)$:
@@
    set all values in the vector $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ to 0.
@@
Next, we construct an example of a secure restaking graph deriving a lower bound on the restaking savings that is linear in the number of services and asymptotically matches to all upper bounds derived so far.
@@
For any $m\in\mathbb{N}$, there are instances of a secure restaking graph $G$ in which the restaking savings $RS(G)\in\Omega(m)$.
@@
Assume the number of validators is $n=m^{2}+1$ for some $m\in\mathbb{N}$. The edge set $E$ consists of $m$ edges $(s,m^{2}+1)$, for any $1\leq s\leq m$ and $m^{2}$ edges of type (s,(s1)m+j)(s,(s-1)m+j), for any $1\leq s\leq m$, $1\leq j\leq m$.
@@
First, we show that $G$ is a secure restaking graph. Note that the large validator, indexed with $m^{2}+1$} and with stake $\sigma_{m^{2}+1}=2m$ does not want to participate in any attack, as she is losing at least the total value available across all services. No service $s\in\{1,2,...,m\}$ can be attacked by all validators securing it other than the large validator, since the service security parameter is $\alpha_{s}=\frac{1}{2m+1}$, while all small validators indexed ((s1)m+1,(s1)k+2,,sm)$($s-1)m+1,(s-1)k+2,...,sm) make up stake $1$ in total. The total available stakes securing a service $s$ is $2m+1$.
@@
    If there is a subset of validators, whose stakes sum up to a number between $1$ and $2$, then it must be that the total stake, $\sum_{i=1}^{m+1}a_{i}$, is at least $2m+1$. Otherwise, this subset would be able to attack the service and the attack would be profitable. This implies that the extra stake size is at least $2m=2m+1-m\frac{1}{m}$.
@@
It is easy to show that $O(m^{2})$ total stake is enough to add to validators in $V$, so that the resulting stake vector $\sigma\in\mathcal{W}(G)$. It can, for example, be done by adding the stake $2m^{2}-m$ to the validator $m^{2}+1$ so that it total stake become $(2m+1)m$. Then splitting this stake into $m$ stakes of $2m+1$ each in each of $m$ PoS protocols. Small stakers in these protocols can not attack any of the services even if they all coordinate, while the large staker does not find any attack profitable. This proves the claim of the theorem, $RS(G)\in\Theta(m)$.
@@
That is, restaking savings are not bounded by a constant, and they may grow unboundedly as the number of services and validators grows. Note that we need both the number of services and the number of validators to grow unboundedly. Otherwise, the upper bounds obtained in Proposition 3 ensure that $RS(G)$ is constant.